OpenStudy (milo123):

MEDAL What is the rational number equivalent to 3.24 with a bar over 24? 3 6/97 3 28/99 3 8/33 3 10/32

OpenStudy (milo123):

@MathGeek289 @Elsa213 @Awolflover1

OpenStudy (mathgeek289):

We Need To Make It A Fraction. We Can Put It Like This.

OpenStudy (mathgeek289):

We Need To Simplify...

rishavraj (rishavraj):

x = 3.24(bar on 24) 100x = 324.24(bar on 24) so on subtracting 99x = 321 find x

OpenStudy (brainzonly):

@rishavraj which according to the golden rule of algebra is dividing 99 from both sides

OpenStudy (mathgeek289):

Never Mind. I Goofed Lol.

rishavraj (rishavraj):

yup @brainzonly :)

OpenStudy (brainzonly):

MathGeek makes everything harder than it is.

OpenStudy (milo123):

LOL

rishavraj (rishavraj):

so x = 321/99 x = 107/33 x = \[3 \frac{ 8 }{ }\]

rishavraj (rishavraj):

oops \[x = 3\frac{ 8 }{ 33 }\]

OpenStudy (brainzonly):

...which is 2.181818181818181818181818181818181818........... forever.

rishavraj (rishavraj):

lmao seriously :P @brainzonly

OpenStudy (milo123):

so the rational number equivenant to 3.242424 = 3 8/33

OpenStudy (brainzonly):

@rishavraj I enjoy stating the obvious. LOL

rishavraj (rishavraj):

yup :)) @milo123 and yeah I can see ..lol @brainzonly

OpenStudy (brainzonly):

I'm waiting for someone to ask a question that's actually hard. :D

OpenStudy (milo123):

ok

OpenStudy (milo123):

http://prnt.sc/cis5ck